# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 3.3. A paraboloidal proof of Słodkowski’s density estimate
The tools above allow one to prove the following “paraboloidal” counterpart of Lemma 3.4 which one finds in [25] which gives a lower bound on the measure of global upper contact points of type $\frac{1}{r}I$ on closed balls as defined in Definition 3.10.
@@
Let $u$ be a convex function on $\overline{B}_{\overline{\rho}}$, such that, for some $R>0$,
$0\leqslant u(y)<\frac{|y|^{2}}{2R}\quad\text{for $y\neq 0$}.$
@@
Apply Proposition 3.19 with $K=\overline{B}_{\rho}$, $\rho\leqslant\overline{\rho}$, and $r\leqslant R$. Since by our assumptions $\operatorname*{osc}_{K}u=M(\rho)\vcentcolon=\sup_{B_{\rho}}u$, we have $\delta=\delta(\rho)=\sqrt{2rM(\rho)}$ and for every $v\in\overline{B}_{\rho-\delta}$, the contact set $\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I,v)$ is a nonempty subset of $B_{\rho}$. Note that, by convexity, if $x$ is a point in the radius $r$ contact set, then the upper vertex map $V_{r}$ is well-defined at $x$. Hence, for every $v\in\overline{B}_{\rho-\delta}$ we have that $v=V_{r}(x)$ for some $x\in\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I,v)\subset\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I)$; that is,
$\overline{B}_{\rho-\delta}\subset V\!\left(\mathcal{C}^{+}(u,\overline{B}_{\rho},\tfrac{1}{r}I)\right).$
@@
$|B_{\rho-\delta}|\leqslant|\mathcal{C}^{+}(u,\overline{B}_{\rho},\tfrac{1}{r}I)|.$ (3.4)
@@
$\rho-\delta>\rho\left(1-\sqrt{\frac{r}{R}}\,\right)\geqslant 0$ (3.5)
@@
On the other hand, a novel “paraboloidal” version of Lemma 3.3 is now easy to formulate and prove.
Lemma 3.21 (Upper bound on the largest eigenvalue of the Hessian: paraboloidal version).
Let $U\subset{\mathbb{R}}^{n}$ be open and $u\colon U\to{\mathbb{R}}$ be convex. Suppose there exists a paraboloid of radius $r$ (i.e., $\overline{\operatorname{epi}}(\varphi)$, with $\varphi$ quadratic of radius $r$), which supports $\Gamma(u)$ from the above at $(x,u(x))$. Then
@@ 3.4. The Jensen–Słodkowski theorem
We begin by using the HL lower bound on the measure of the set of lower contact points in Lemma 3.20 to prove the HL version of the Słodkowski lemma as done in [25].
@@
Furthermore, note that the HL–Słodkowski lemma can be paraphrased by saying that if $u$ is locally convex and its quadratic perturbation $u-\frac{\lambda}{2}|\cdot-\,x|^{2}$ has a strict local maximum at $x$, then the set of global upper contact points of type $\lambda I$ for $u$ on each small ball about $x$ has positive measure. Analogously, Jensen’s Lemma 3.6 considers a quasi-convex function with a strict local maximum at $\hat{x}$ and states that the set of the upper contact points near $\hat{x}$ whose associated upper contact jets are of the form $(p,0)$, with $p\in{\mathbb{R}}^{n}$ small ($|p|\leqslant\delta$), has positive measure.
With the intuition that a bridge between convex and quasi-convex functions can relate the two lemmas, Harvey and Lawson [25] prove that the HL–Słodkowski lemma is equivalent to the following result, which they simply called Jensen’s lemma. It roughly states that if $u$ is locally quasi-convex, then the set of contact points of locally supporting hyperplanes from above near a strict local maximum point cannot be too small.
@@
Suppose that $w$ is a locally quasi-convex function on $X\subset{\mathbb{R}}^{n}$ with strict upper contact jet $(0,0)$ at $x$. Then there exists $\overline{\rho}>0$ such that
@@
Consider $u$ and $w$ related by the identity $u=w+\frac{\lambda}{2}|\cdot-\,x|^{2}$. Then, on $B_{\rho}(x)$, $u$ is convex if and only if $w$ is $\lambda$-quasi-convex; note that by Definition 2.1, the parameter of quasi-convexity $\lambda(x)$ is locally constant, therefore we may assume that the locally quasi-convex function $w$ is $\lambda$-quasi-convex on $B_{\rho}(x)$ if $\rho>0$ small enough. Furthermore, $(0,\lambda I)$ is a strict upper contact jet for $u$ at $x$ if and only if $(0,0)$ is for $w$. This shows that $\mathcal{C}^{+}(w,\overline{B}_{\rho}(x),0)=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),\lambda I)$, and completes the proof of the desired equivalence. ∎
@@
Note that the request that $w$ has a strict local maximum at $\hat{x}$ is equivalent to the request that $(0,0)$ is a strict upper contact jet for $w$ at $\hat{x}$, which for $w$ $\lambda$-quasi-convex is equivalent to the requirement that $(0,\lambda I)$ is a strict upper contact jet for $u\vcentcolon=w+\frac{\lambda}{2}|\cdot-\,\hat{x}|^{2}$ at $\hat{x}$. Furthermore, $w_{p}$ has a local maximum at $x$ if and only if $(-p,0)$ is an upper contact jet for $w$ at $x$; hence
@@
where the set $\mathcal{C}^{+}(w,\overline{B}_{r}(\hat{x}),0\,;\,p)$ denotes the set of the global upper contact points on $\overline{B}_{r}(\hat{x})$ for $w$ associated to the jet $(p,0)$. It is easy to see that $\mathcal{C}^{+}(w,\overline{B}_{r}(\hat{x}),0\,;\,p)$ is the set of those points $x\in\overline{B}_{r}(\hat{x})$ such that $(p+\lambda(x-\hat{x}),\lambda I)$ is an upper global upper contact jet (of radius $r=1/\lambda$) for $u$ at $x$. By Lemma 2.13, $Du(x)=p+\lambda(x-\hat{x})$ and thus
$V_{\frac{1}{\lambda}}(x):=\left(I-\frac{1}{\lambda}Du\right)(x)=\hat{x}-\frac{p}{\lambda}\quad\implies\quad|p|=\lambda\,|V_{\frac{1}{\lambda}}(x)-\hat{x}|,$
where $V_{\frac{1}{\lambda}}\colon\mathcal{C}^{+}(u,\overline{B}_{r}(\hat{x}),\lambda I)\to{\mathbb{R}}^{n}$ is the upper vertex map of radius $r=1/\lambda$ associated to $u$. Since $V_{\frac{1}{\lambda}}$ is nonexpansive (Proposition 3.18) and $V_{\frac{1}{\lambda}}(\hat{x})=\hat{x}$, we see that
@@
We now merge the two equivalent Lemmas 3.22 and 3.23 into a theorem which, despite being another equivalent reformulation of the previous results, unveils their full generality. This is also done in [25].
@@
Suppose that $w$ is a locally quasi-convex function with strict upper contact jet $(p,A)$ at $x$. Then there exists $\overline{\rho}>0$ such that
@@
for some $\lambda>0$ to be determined. It is easy to see that $(0,\lambda I)$ is a strict upper contact jet for $u\vcentcolon=w+\varphi$ at $x$, and $\mathcal{C}^{+}(w,\overline{B}_{\rho}(x),A)=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),\lambda I)$. Moreover, let $\alpha,\beta>0$ such that $w$ is $\alpha$-quasi-convex and $A\leqslant\beta I$; then $u$ is convex if $\lambda\geqslant\alpha+\beta$. The thesis now follows from Lemma 3.22. ∎
@@
3.5. Consequences of the Jensen-Słodkowski Theorem
As previously mentioned, the main consequence of the Jensen–Słodkowski theorem (Theorem 3.28) that concerns us in this work is the property of partial upper semicontinuity of second derivatives of a locally quasi-convex function (PUSC of SD, Lemma 2.15). We will now give the proof, which also competes the proof of the Upper Contact Jet Theorem (Theorem 2.12).
@@ 3.5. Consequences of the Jensen-Słodkowski Theorem
Note that this is certainly possible for every sufficiently small $\varepsilon>0$ since $E\cap\operatorname{Diff}^{2}(u)$ has full measure in a neighborhood of $x$ by Alexandrov’s Theorem 2.2 and $(p,A+\varepsilon I)$ is a strict upper contact jet for $u$ at $x$, therefore $\mathcal{C}^{+}(u,\overline{B}_{\varepsilon}(x),A+\varepsilon I)$ has positive measure for $\varepsilon$ small by the Jensen–Słodkowski Theorem 3.28. We know that since $u$ is twice differentiable at $x_{k}$, then $D^{2}u(x_{k})\leqslant A+\varepsilon_{k}I$. Furthermore, by quasi-convexity, $D^{2}u(x_{k})+\lambda I\geqslant 0$ for some $\lambda\geqslant 0$. By compactness, there exists a subsequence $\{x_{j}\}$ such that $D^{2}u(x_{j})\to\bar{A}$ for some $\bar{A}\in[-\lambda I,A]$. ∎
We also wish to note here that, as the HL–Słodkowski and the HL–Jensen lemmas are equivalent, despite dealing with convex and quasi-convex functions, respectively, one should expect results coming from the former lemma to have a counterpart holding for quasi-convex functions as well. This is the case, for instance, of Słodkowski’s Largest Eigenvalue Theorem 3.5. In fact, if $u$ is $\lambda$-quasi-convex near $x^{*}$ with $\mathcal{K}(u,x)=k^{*}$, then $\tilde{u}\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ is convex with $\mathcal{K}(\tilde{u},x^{*})=\mathcal{K}(u,x^{*})+\lambda=k^{*}+\lambda$. Then Lemma 3.1 tells that, for $k>k^{*}$,
@@
Exploiting this density estimate for quasi-convex functions one can deduce a quasi-convex version of Słodkowski’s Largest Eigenvalue Theorem. Alternatively, it is possible to prove it directly from the Jensen–Słodkowski Theorem 3.28, as we do here.
Theorem 3.30 (Słodkowski’s Largest Eigenvalue Theorem: quasi-convex version).
Let $u$ be locally quasi-convex on an open set $X$ and suppose $K(u,x)\geqslant M$ for a.e. $x\in X$. Then $K(u,x)\geqslant M$ for all $x\in X$.
@@
Suppose that $u$ is locally quasi-convex on $X$ open and that $\mathcal{K}(u,x)\geqslant M$ for a.e. $x\in X$. Seeking a contradiction to $\mathcal{K}(u,x)\geqslant M$ for every $x\in X$, suppose that there exists $\hat{x}\in X$ such that $\mathcal{K}(u,\hat{x})<M$; say $\mathcal{K}(u,\hat{x})=M-3\delta$ for some $\delta>0$. By the definition (3.1) of $\mathcal{K}$, there exists $\overline{\varepsilon}>0$ such that $\overline{B}_{\overline{\varepsilon}}(\hat{x})\subset X$ and
@@
This last inequality is equivalent to $(Du(\hat{x}),(M-2\delta)I)$ being an upper contact jet for $u$ at $\hat{x}$. Hence, $(Du(\hat{x}),(M-\delta)I)$ is strict and by Theorem 3.28 we have
@@ 3.6. An area formula proof of the Jensen–Słodkowski theorem
We conclude this section with the very interesting fact observed by Reese Harvey [20]; namely, that the Jensen–Słodkowski theorem can be seen as an immediate consequence of Alexandrov’s maximum principle, a short proof of which can be in turn built around the well-known area formula (see [19, Area Theorem 3.2.3]). Therefore, if one also considers that the Lipschitz version of Sard’s theorem we used in the proof of Alexandrov’s theorem can be seen as a corollary of the area formula as well (cf. [13, Notes on the appendix]), [^4] The proof of Sard’s theorem we propose in Appendix LABEL:proofsard, based on Besicovitch’s covering theorem, could seem at first sight more involuted than just invoking the area formula, yet the reader should be aware that classical proofs of the area formula itself (see, e.g. [18]) rely, among other instruments, on the Radon–Nikodym theorem, which in turn also requires Besicovitch’s result. then the Area Theorem arises as the pillar of the theory we discussed so far.
We recall the following definitions of approximate limit superior and approximate differential (for further details, see [19, p. 212]).
@@
Let $g$ be a measurable function in a neighborhood of a point $x$. The approximate limit superior of $g$ at $x$ is the quantity
$\operatorname*{ap\,lim\,sup}_{y\to x}g(y)\vcentcolon=\inf\left\{t\in{\mathbb{R}}\ :\ \lim_{r\searrow 0}\frac{|g^{-1}((t,+\infty))\cap B_{r}(x)|}{|B_{r}(x)|}=0\right\}.$
@@
Now, with $f$ be as in Definition 3.32 if one defines the Stepanov set of $f$ by
@@
then by H. Federer’s extension of W. Stepanoff’s theorem [19, Lemma 3.1.7 and Theorem 3.1.8], $f$ has an approximate differential almost everywhere in $S(f)$; also, as noted in the beginning of [19, Section 3.2], the Area Theorem can be stated as follows.
@@
Let $u\colon X\to{\mathbb{R}}$ be locally quasi-convex on $X$ open in ${\mathbb{R}}^{n}$ and consider the gradient map $Du\colon\operatorname{Diff}^{1}(u)\to{\mathbb{R}}^{n}$. We know that both $\operatorname{Diff}^{1}(u)$ and $\operatorname{Diff}^{2}(u)$ have full measure in $X$, and hence the next result follows easily.
@@
Theorem 3.35 (Area Theorem for gradients of locally quasi-convex functions).
Given $u:X\to{\mathbb{R}}$ a locally quasi-convex function. For every measurable set $E\subset\operatorname{Diff}^{2}(u)$, one has
$|\Sigma|:=|Du(E)|\leqslant\int_{E}|\det D^{2}u(x)|\,\mathrm{d}x.$ (3.6)
Next, we use this Area Theorem 3.35 for locally quasi-convex functions to prove a locally quasi-convex version of Alexandrov’s maximum principle, which takes the form of the following pointwise a priori estimate.
Theorem 3.36 (Alexandrov’s maximum principle for locally quasi-convex functions).
Let $\Omega\subset{\mathbb{R}}^{n}$ be open and bounded. If $u\in\operatorname{USC}(\overline{\Omega})$ is locally quasi-convex, then
$\sup_{\overline{\Omega}}u\leqslant\sup_{\partial\Omega}u^{+}+\frac{\mathrm{diam}(\Omega)}{\omega_{n}^{1/n}}\Bigg{(}\int_{E}|\det D^{2}u(x)|\,\mathrm{d}x\Bigg{)}^{1/n},$ (3.7)
where $\omega_{n}:=|B_{1}|$ is the Lebesgue measure of the unit ball in ${\mathbb{R}}^{n}$ and
$E:=\mathcal{C}^{+}(u,\Omega,0)\cap\operatorname{Diff}^{2}(u)$ (3.8)
is the set of all flat global upper contact points at which $u$ is twice differentiable.
@@
The idea of the proof is to show that the image of the gradient map $\Sigma:=Du(E)$ (or some superset of $\Sigma$ with the same measure) contains an open ball $B_{\delta}$ (about the origin) of a suitable radius $\delta$. In particular, if one defines
@@
Now, if one can show that $B_{\delta}\subset\Sigma:=Du(E)$, then

### v3

@@ 3.3. A paraboloidal proof of Słodkowski’s density estimate
The tools above allow one to prove the following “paraboloidal” counterpart of Lemma 3.4 which one finds in [28] and which gives a lower bound on the measure of global upper contact points of type $\frac{1}{r}I$ on closed balls as defined in Definition 3.10.
@@
Let $u$ be a convex function on $\overline{B}_{\overline{\rho}}$ such that, for some $R>0$,
$0\leqslant u(y)<\frac{|y|^{2}}{2R}\quad\text{if $y\neq 0$}.$
@@
Apply Proposition 3.18 with $K=\overline{B}_{\rho}$, $\rho\leqslant\overline{\rho}$, and $r\leqslant R$. Since by our assumptions $\operatorname*{osc}_{K}u=M(\rho)\vcentcolon=\sup_{B_{\rho}}u$, we have $\delta=\delta(\rho)=\sqrt{2rM(\rho)}$ and for every $v\in\overline{B}_{\rho-\delta}$, the contact set $\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I,v)$ is a nonempty subset of $B_{\rho}$. Note that, by convexity, if $x$ is a point in the contact set of type $\frac{1}{r}I$, then the upper vertex map $V_{r}$ is well-defined at $x$. Hence, for every $v\in\overline{B}_{\rho-\delta}$ we have that $v=V_{r}(x)$ for some $x\in\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I,v)\subset\mathcal{C}^{+}(u,\overline{B}_{\rho},\frac{1}{r}I)$; that is,
$\overline{B}_{\rho-\delta}\subset V_{r}\!\left(\mathcal{C}^{+}(u,\overline{B}_{\rho},\tfrac{1}{r}I)\right).$
@@
$|B_{\rho-\delta}|\leqslant|\mathcal{C}^{+}(u,\overline{B}_{\rho},\tfrac{1}{r}I)|.$
@@
$\rho-\delta>\rho\left(1-\sqrt{\frac{r}{R}}\,\right)\geqslant 0$
@@
A novel paraboloidal version of Lemma 3.3 is now easy to formulate and prove.
Lemma 3.20 (Upper bound on the largest eigenvalue of the Hessian; paraboloidal version).
Let $u\colon U\to{\mathbb{R}}$ be convex on $U\subset{\mathbb{R}}^{n}$ open and convex. Suppose that there exists a paraboloid of radius $r$ which supports $\Gamma(u)$ from above at $(x,u(x))$; that is, there exits a quadratic function $\varphi$ of radius $r$ such that $\operatorname{epi}(\varphi)$ lies above the graph $\Gamma(u)$ of $u$ and touches $\Gamma(u)$ at $(x,u(x))$. Then one has
@@ 3.4. The Jensen–Słodkowski theorem
We begin by using the lower bound on the measure of the set of lower contact points in Lemma 3.19 to prove the HL version of the Słodkowski lemma as done in [28].
@@
Furthermore, note that the HL–Słodkowski lemma can be paraphrased by saying that if $u$ is locally convex and its quadratic perturbation $u-\frac{\lambda}{2}|\cdot-\,x|^{2}$ has a strict local maximum at $x$, then the set of global upper contact points of type $\lambda I$ for $u$ on each small ball about $x$ has positive measure. Analogously, Jensen’s Lemma 3.6 considers a semiconvex function with a strict local maximum at $\hat{x}$ and states that the set of the upper contact points near $\hat{x}$ whose associated upper contact jets are of the form $(p,0)$, with $p\in{\mathbb{R}}^{n}$ small ($|p|\leqslant\delta$), has positive measure.
With the intuition that a bridge between convex and semiconvex functions can relate the two lemmas, Harvey and Lawson [28] prove that the HL–Słodkowski lemma is equivalent to the following result, which they simply call Jensen’s lemma. It roughly states that if $u$ is locally semiconvex, then the set of contact points of locally supporting hyperplanes from above near a strict local maximum point cannot be too small.
@@
Suppose that $w$ is a locally semiconvex function on $X\subset{\mathbb{R}}^{n}$ with strict upper contact jet $(0,0)$ at $x$. Then there exists $\overline{\rho}>0$ such that
@@
Consider $u$ and $w$ related by the identity $u=w+\frac{\lambda}{2}|\cdot-\,x|^{2}$. Then, on $B_{\rho}(x)$, $u$ is convex if and only if $w$ is $\lambda$-semiconvex; note that by Definition 2.1, the parameter of semiconvexity $\lambda(x)$ is locally constant, therefore we may assume that the locally semiconvex function $w$ is $\lambda$-semiconvex on $B_{\rho}(x)$ if $\rho>0$ small enough. Furthermore, $(0,\lambda I)$ is a strict upper contact jet for $u$ at $x$ if and only if $(0,0)$ is for $w$. This shows that $\mathcal{C}^{+}(w,\overline{B}_{\rho}(x),0)=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),\lambda I)$, and completes the proof of the desired equivalence. ∎
@@
Note that the request that $w$ has a strict local maximum at $\hat{x}$ is equivalent to the request that $(0,0)$ is a strict upper contact jet for $w$ at $\hat{x}$, which for $w$ $\lambda$-semiconvex is equivalent to the requirement that $(0,\lambda I)$ is a strict upper contact jet for $u\vcentcolon=w+\frac{\lambda}{2}|\cdot-\,\hat{x}|^{2}$ at $\hat{x}$. Furthermore, $w_{p}$ has a local maximum at $x$ if and only if $(-p,0)$ is an upper contact jet for $w$ at $x$; hence
@@
where the set $\mathcal{C}^{+}(w,\overline{B}_{r}(\hat{x}),0\,;\,p)$ denotes the set of the global upper contact points on $\overline{B}_{r}(\hat{x})$ for $w$ associated to the jet $(p,0)$. It is easy to see that $\mathcal{C}^{+}(w,\overline{B}_{r}(\hat{x}),0\,;\,p)$ is the set of those points $x\in\overline{B}_{r}(\hat{x})$ such that $(p+\lambda(x-\hat{x}),\lambda I)$ is a global upper contact jet for $u$ at $x$. By Lemma 2.13, $Du(x)=p+\lambda(x-\hat{x})$ and thus
$V_{\frac{1}{\lambda}}(x)\vcentcolon=\Bigl{(}I-\frac{1}{\lambda}Du\Bigr{)}(x)=\hat{x}-\frac{p}{\lambda}\quad\implies\quad|p|=\lambda\,|V_{\frac{1}{\lambda}}(x)-\hat{x}|,$
where $V_{\frac{1}{\lambda}}\colon\mathcal{C}^{+}(u,\overline{B}_{r}(\hat{x}),\lambda I)\to{\mathbb{R}}^{n}$ is the upper vertex map of radius $r=\frac{1}{\lambda}$ associated to $u$. Since $V_{\frac{1}{\lambda}}$ is nonexpansive (Proposition 3.17) and $V_{\frac{1}{\lambda}}(\hat{x})=\hat{x}$, we see that
@@
We now merge the two equivalent Lemmas 3.21 and 3.22 into a theorem which, despite being another equivalent reformulation of the previous results, unveils their full generality. This is also done in [28].
@@
Suppose that $w$ is a locally semiconvex function with strict upper contact jet $(p,A)$ at $x$. Then there exists $\overline{\rho}>0$ such that
@@
for some $\lambda>0$ to be determined. It is easy to see that $(0,\lambda I)$ is a strict upper contact jet for $u\vcentcolon=w+\varphi$ at $x$, and $\mathcal{C}^{+}(w,\overline{B}_{\rho}(x),A)=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),\lambda I)$. Moreover, let $\alpha,\beta>0$ such that $w$ is $\alpha$-semiconvex and $A\leqslant\beta I$; then $u$ is convex if $\lambda\geqslant\alpha+\beta$. The thesis now follows from Lemma 3.21. ∎
@@
3.5. Consequences of the Jensen–Słodkowski Theorem
As previously mentioned, the main consequence of the Jensen–Słodkowski Theorem 3.27 that concerns us in this work is the property of partial upper semicontinuity of second derivatives of a locally semiconvex function stated in Lemma 2.15(PUSC of SD). We will now give the proof, which also competes the proof of the Upper Contact Jet Theorem 2.12.
@@ 3.5. Consequences of the Jensen–Słodkowski Theorem
Note that this is certainly possible for every sufficiently small $\varepsilon>0$ since $E\cap\operatorname{Diff}^{2}(u)$ has full measure in a neighborhood of $x$ by Alexandrov’s Theorem 2.2 and $(p,A+\varepsilon I)$ is a strict upper contact jet for $u$ at $x$, therefore $\mathcal{C}^{+}(u,\overline{B}_{\varepsilon}(x),A+\varepsilon I)$ has positive measure for $\varepsilon$ small by the Jensen–Słodkowski Theorem 3.27. We know that since $u$ is twice differentiable at $x_{k}$, then $D^{2}u(x_{k})\leqslant A+\varepsilon_{k}I$. Furthermore, by semiconvexity, $D^{2}u(x_{k})+\lambda I\geqslant 0$ for some $\lambda\geqslant 0$. By compactness, there exists a subsequence $\{x_{j}\}$ such that $D^{2}u(x_{j})\to\bar{A}$ for some $\bar{A}\in[-\lambda I,A]$. ∎
We also wish to note that, as the HL–Słodkowski and the HL–Jensen lemmas are equivalent, despite dealing with convex and semiconvex functions, respectively, one should expect results coming from the former lemma to have a counterpart holding for semiconvex functions as well. This is the case, for instance, of Słodkowski’s Largest Eigenvalue Theorem 3.5. In fact, if $u$ is $\lambda$-semiconvex near $x^{*}$ with $\mathcal{K}(u,x)=k^{*}$, then $\tilde{u}\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ is convex with $\mathcal{K}(\tilde{u},x^{*})=\mathcal{K}(u,x^{*})+\lambda=k^{*}+\lambda$. Then Lemma 3.2 tells that, for $k>k^{*}$,
@@
Exploiting this density estimate for semiconvex functions one can deduce a semiconvex version of Słodkowski’s Largest Eigenvalue Theorem. Alternatively, it is possible to prove it directly from the Jensen–Słodkowski Theorem 3.27, as we do here.
Theorem 3.29 (Słodkowski’s Largest Eigenvalue Theorem: semiconvex version).
Let $u$ be locally semiconvex on an open set $X$ and suppose $\mathcal{K}(u,x)\geqslant M$ for a.e. $x\in X$. Then $\mathcal{K}(u,x)\geqslant M$ for all $x\in X$.
@@
Suppose that $u$ is locally semiconvex on $X$ open and that $\mathcal{K}(u,x)\geqslant M$ for a.e. $x\in X$. Seeking a contradiction to $\mathcal{K}(u,x)\geqslant M$ for every $x\in X$, suppose that there exists $\hat{x}\in X$ such that $\mathcal{K}(u,\hat{x})<M$; say $\mathcal{K}(u,\hat{x})=M-3\delta$ for some $\delta>0$. By the definition (3.1) of $\mathcal{K}$, there exists $\overline{\varepsilon}>0$ such that $\overline{B}_{\overline{\varepsilon}}(\hat{x})\subset X$ and
@@
This last inequality is equivalent to $(Du(\hat{x}),(M-2\delta)I)$ being an upper contact jet for $u$ at $\hat{x}$. Hence, $(Du(\hat{x}),(M-\delta)I)$ is strict and by the Jensen–Słodkowski  Theorem 3.27 we have
@@ 3.6. An area formula proof of the Jensen–Słodkowski theorem
We conclude this chapter with a very interesting fact observed by Reese Harvey [23]; namely, that the Jensen–Słodkowski theorem can be seen as an immediate consequence of Alexandrov’s maximum principle, a short proof of which can be in turn built around the well-known area formula (see [20, Area Theorem 3.2.3]). Therefore, if one also considers that the Lipschitz version of Sard’s theorem used in the proof of Alexandrov’s theorem can be seen as a corollary of the area formula as well (cf. [14, Notes on the appendix]), [^1] The proof of Sard’s theorem we propose in Appendix B, based on Besicovitch’s covering theorem, could seem at first sight more involuted than just invoking the area formula, yet the reader should be aware that classical proofs of the area formula itself (see, e.g. [19]) rely, among other instruments, on the Radon–Nikodym theorem, which in turn also requires Besicovitch’s result. then the Area Theorem arises as the pillar of the theory we discussed so far.
We recall the following definitions of approximate (upper/lower) limit and approximate differential; for further details, see [20, p. 212].
@@
Let $g$ be a measurable function in a neighborhood of a point $x$. The approximate upper limit of $g$ at $x$ is the quantity
$\operatorname*{ap\,lim\,sup}_{y\to x}g(y)\vcentcolon=\inf\left\{t\in{\mathbb{R}}\ :\ \lim_{r\searrow 0}\frac{|g^{-1}((t,+\infty))\cap B_{r}(x)|}{|B_{r}(x)|}=0\right\};$
symmetrically, the approximate lower limit is
$\operatorname*{ap\,lim\,inf}_{y\to x}g(y)\vcentcolon=\sup\left\{t\in{\mathbb{R}}\ :\ \lim_{r\searrow 0}\frac{|g^{-1}((-\infty,t)\cap B_{r}(x)|}{|B_{r}(x)|}=0\right\}.$
Then, the approximate limit, $\operatorname*{ap\,lim}_{y\to x}g(y)$, is the common value of approximate upper and lower limits, if it exists.
@@
If, for $f$ as in Definition 3.31, one defines the Stepanov set of $f$ by
@@
then by Federer’s extension of Stepanov’s theorem [20, Lemma 3.1.7 and Theorem 3.1.8], $f$ has an approximate differential almost everywhere in $S(f)$; also, as noted in the beginning of [20, Section 3.2], the Area Theorem can be stated as follows.
@@
Let $u\colon X\to{\mathbb{R}}$ be locally semiconvex on $X$ open in ${\mathbb{R}}^{n}$ and consider the gradient map $Du\colon\operatorname{Diff}^{1}(u)\to{\mathbb{R}}^{n}$. We know that both $\operatorname{Diff}^{1}(u)$ and $\operatorname{Diff}^{2}(u)$ have full measure in $X$, and hence the next result follows easily.
@@
Theorem 3.34 (Area Theorem for gradients of locally semiconvex functions).
Given $u\colon X\to{\mathbb{R}}$ a locally semiconvex function, for every measurable set $E\subset\operatorname{Diff}^{2}(u)$ one has
$|Du(E)|\leqslant\int_{E}|\det D^{2}u(x)|\,\mathrm{d}x.$ (3.4)
Next, we use this Area Theorem 3.34 for locally semiconvex functions to prove a locally semiconvex version of Alexandrov’s maximum principle, which takes the form of the following pointwise a priori estimate.
Theorem 3.35 (Alexandrov’s maximum principle for locally semiconvex functions).
Let $\Omega\subset{\mathbb{R}}^{n}$ be open and bounded. If $u\in\operatorname{USC}(\overline{\Omega})$ is locally semiconvex, then
$\sup_{\overline{\Omega}}u\leqslant\sup_{\partial\Omega}u+\frac{\mathrm{diam}(\Omega)}{\omega_{n}^{1/n}}\biggl{(}\int_{E(u)}|\det D^{2}u(x)|\,\mathrm{d}x\biggr{)}^{1/n},$ (3.5)
where $\omega_{n}\vcentcolon=|B_{1}|$ is the Lebesgue measure of the unit ball in ${\mathbb{R}}^{n}$ and $E(u)$ is the set of all flat global upper contact points at which $u$ is twice differentiable; that is,
$E(u)\vcentcolon=\mathcal{C}^{+}(u,\Omega,0)\cap\operatorname{Diff}^{2}(u).$ (3.6)
@@
It suffices to prove the weaker inequality [^2] It is weaker since $\sup_{\partial\Omega}u\leqslant\sup_{\partial\Omega}u^{+}$.
$\sup_{\overline{\Omega}}u\leqslant\sup_{\partial\Omega}u^{+}+\frac{\mathrm{diam}(\Omega)}{\omega_{n}^{1/n}}\biggl{(}\int_{E(u)}|\det D^{2}u(x)|\,\mathrm{d}x\biggr{)}^{1/n}.$ (3.7)
Indeed, the translated function $\bar{u}\vcentcolon=u-\sup_{\partial\Omega}u$, where $\sup_{\partial\Omega}u<+\infty$ since $u\in\operatorname{USC}(\overline{\Omega})$, also satisfies $\bar{u}\in\operatorname{USC}(\overline{\Omega})$ and $\bar{u}$ is locally semiconvex. Moreover, one clearly has
$\bar{u}^{+}=0\ \text{on}\ \partial\Omega\quad\text{and}\quad D^{2}\bar{u}=D^{2}u\ \text{on}\ E\vcentcolon=E(\bar{u})=E(u).$ (3.8)
Applying the weaker inequality (3.7) to $\bar{u}$ yields
[…]
which, using the definition of $\bar{u}$ and the properties (3.8), is equivalent to
[…]
and hence (3.5) for $u$.
The idea of the proof of (3.7) is to show that the image of the gradient map $Du(E)$ (or some superset of it with the same measure) contains an open ball $B_{\delta}$ of a suitable radius $\delta$; in particular, it suffices to choose
@@
Now, if one can show that $B_{\delta}\subset Du(E)$, then